40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 17 00:55:30 2010, Koji, Update, IOO, The dark port beams reached the AP table 
    Reply  Sat Dec 18 01:17:04 2010, kiwamu, Update, IOO, some more alignments 
Message ID: 4070     Entry time: Sat Dec 18 01:17:04 2010     In reply to: 4067
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: some more alignments 

 [Koji and Kiwamu]

 We did some more vacuum works today. It is getting ready to pump down.

 (what we did) 

 - alignment of the POY mirrors. Now the beam is coming out from the ITMY chamber successfully

  - leveling of the tables (except for the IOO and OMC chamber)

  - realigned the beam axis down to the Y arm because the leveling of the BS table changed the alignments.

 - installed IP_POS mirrors

 - aligned the green beam and made it overlap with IR beam path.

 - repositioned green steering mirrors since one of them are too close to the dark beam path

 - aligned everything.

ELOG V3.1.3-